OpenStudy (anonymous):

Solve -3(2x - 4) = 24 if anyone could teach me the process to solve this

OpenStudy (anonymous):

distribute

OpenStudy (aripotta):

ok, first, use the distributive property to distribute the -3

OpenStudy (anonymous):

so its -3 * 2x then -3 * -4

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so -6x + 12 = 24

OpenStudy (anonymous):

-6x = 12 24- 12 = 12

OpenStudy (anonymous):

divide both sides by -6 and x = ?

OpenStudy (anonymous):

12/-6 = ?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

my bad a bit lat eon the response but thank you!

OpenStudy (anonymous):

-3(2x - 4) = 24 1st Distributive property so, -3(2x) -3(-4) = 24 substract 12 from each side -6x + 12 = 24 combine like terms -6x +12 - 12 = 24 -12 divide both side by -6 -6x = 12 -6x/ -6 = 12/-6 x = -2